OpenStudy (anonymous):

Setup and solve the following: Find two positive real numbers that differ by 2 and have a product of 10 @Mathematics

OpenStudy (anonymous):

x-y=2 xy=10

OpenStudy (anonymous):

y=x-2

OpenStudy (anonymous):

x(x-2)=10

OpenStudy (anonymous):

x^2-2x=10

OpenStudy (anonymous):

its a completeing the square problem?

OpenStudy (anonymous):

no quadratic formula

OpenStudy (anonymous):

Oh I guess you could do either